### v1

# Learning What Reinforcement Learning Can’t: Interleaved Online Fine-Tuning for hardest Questions

Abstract
Recent advances in large language model (LLM) reasoning have shown that sophisticated behaviors such as planning and self-reflection can emerge through reinforcement learning (RL). However, despite these successes, RL in its current form remains insufficient to induce capabilities that exceed the limitations of the base model, as it is primarily optimized based on existing knowledge of the model rather than facilitating the acquisition of new information. To address this limitation, we employ supervised fine-tuning (SFT) to learn what RL cannot, which enables the incorporation of new knowledge and reasoning patterns by leveraging high-quality demonstration data. We analyze the training dynamics of RL and SFT for LLM reasoning and find that RL excels at maintaining and improving performance on questions within the model’s original capabilities, while SFT is more effective at enabling progress on questions beyond the current scope of the model. Motivated by the complementary strengths of RL and SFT, we introduce a novel training approach, ReLIFT (Reinforcement Learning Interleaved with Online Fine-Tuning). In ReLIFT, the model is primarily trained using RL, but when it encounters challenging questions, high-quality solutions are collected for fine-tuning, and the training process alternates between RL and fine-tuning to enhance the model’s reasoning abilities. ReLIFT achieves an average improvement of over +5.2 points across five competition-level benchmarks and one out-of-distribution benchmark compared to other zero-RL models. Furthermore, we demonstrate that ReLIFT outperforms both RL and SFT while using only 13% of the detailed demonstration data, highlighting its scalability. These results provide compelling evidence that ReLIFT overcomes the fundamental limitations of RL and underscores the significant potential.
Project Page: https://github.com/TheRoadQaQ/ReLIFT
1 Introduction
Recent advancements in large language model (LLM) reasoning, as demonstrated by OpenAI-o1 Jaech et al. (2024), DeepSeek-R1 Guo et al. (2025), and Kimi-1.5 Team et al. (2025), highlight the potential to enhance reasoning capabilities by increasing test-time computational resources. These models typically generate long Chains-of-Thought (CoT Wei et al. (2022)) responses, exhibiting sophisticated behaviors such as reflection Kumar et al. (2024); Team et al. (2025) and planning Team et al. (2025). The primary factor driving this progress is large-scale Reinforcement Learning with Verifiable Rewards (RLVR), which does not depend on traditional scaffolding techniques like Monte Carlo Tree Search Xie et al. (2024); Li et al. (2025) or Process Reward Models Setlur et al. (2024). In RLVR, rewards are assigned based on whether the model’s output matches a ground-truth solution in mathematics or passes unit tests in code, thereby enabling scalability without the need for detailed demonstration data. In practice, RLVR is typically implemented using policy optimization algorithms such as Proximal Policy Optimization (PPO) Schulman et al. (2017) and Group Relative Policy Optimization (GRPO) Shao et al. (2024). Large-scale RLVR, characterized by its simplicity and directness, has enabled the emergence of long CoT reasoning and self-reflection capabilities.
However, despite the empirical successes of RLVR, its current form is insufficient to incentivize capabilities that transcend the base model’s limitations Yue et al. (2025); Zhao et al. (2025). RL is inherently on-policy, learning from the LLM’s self-generated responses through iterative rollouts and feedbacks. Essentially, RL primarily enhances performance by biasing the model toward reasoning paths it already knows are more likely to yield rewards. In other words, most RL methods optimize based on the model’s existing knowledge rather than facilitating the acquisition of new information. In contrast, supervised fine-tuning (SFT) offers a complementary approach to enhancing the reasoning capabilities of LLMs. By leveraging high-quality demonstration data, SFT facilitates the incorporation of new knowledge and reasoning patterns Mecklenburg et al. (2024); Yue et al. (2025), and has been shown to achieve superior reasoning performance compared to RL, particularly in small-scale models Guo et al. (2025). Nevertheless, the effectiveness of SFT is highly contingent upon the availability of substantial amounts of high-quality demonstration data. Furthermore, models trained with SFT often exhibit limited generalization to out-of-distribution (OOD) scenarios, whereas RL-based approaches tend to demonstrate greater robustness Chu et al. (2025). These contrasting characteristics of RL and SFT highlight a promising research direction:
How can RL and SFT be effectively combined to improve LLM reasoning and OOD generalization, reduce dependence on expensive demonstration data, and achieve capabilities beyond current cognitive constraints?
To achieve this goal, we analyze the training dynamics by examining how the accuracy of the questions change under both RL and SFT. In Section 2.1, we evaluate the accuracy of the model on the questions at various checkpoints and classify these questions into four levels of difficulty. Our findings show that RL is more effective for relatively easy questions, whereas SFT is more beneficial for most difficult ones. Furthermore, for simple questions, SFT can degrade the model’s existing performance, as accuracy on these questions decreases and response length increases. In contrast, for challenging questions, the improvement provided by RL is less pronounced compared to SFT. Overall, RL efficiently trains the model to generate correct answers for questions it can already solve, while SFT is crucial for enabling the model to address questions that surpass its current capabilities.
Based on these analyses, we propose Reinforcement Learning Interleaved with Online Fine-Tuning (ReLIFT). As illustrated in Figure 2, during RL training, we collect highly challenging examples according to the accuracy observed during the rollout. When such a challenging question is identified, we obtain a high-quality CoT solution and subsequently filter out any incorrect answers. After constructing these examples, we add them to an SFT buffer. Once the number of challenging questions in the SFT buffer is sufficient, we perform SFT on these questions for one step.
We conduct comprehensive experiments on five competition-level math reasoning benchmarks and one out-of-distribution (OOD) benchmark. Our proposed ReLIFT method achieves a new state-of-the-art accuracy of 51.1% based on Qwen2.5-Math-7B, outperforming previous zero-RL baselines by a significant margin, surpassing the strongest zero RL baseline by +5.2 points. Notably, ReLIFT demonstrates clear advantages over both pure SFT and pure RL, while requiring only 13% amount of detailed demonstration data. In addition, ReLIFT generates much more concise solutions, reducing average response length by nearly 10x compared to SFT, thus improving both performance and efficiency. We further validate the generality of ReLIFT by extending it to smaller and weaker foundation models, consistently observing superior results over standard RL and SFT approaches.
2 Reinforcement Learning Interleaved with Online Fine-Tuning
2.1 RL vs. SFT: Training Dynamics Across Question Difficulty Levels
We start our study by examining how the model’s accuracy evolves on questions of varying difficulty during both SFT and RL training. Specifically, we independently perform SFT and RL on the Qwen2.5-Math-7B model, using a subset of the Open-R1-math-220K Face (2025). Both SFT and RL are conducted for 120 steps, with model checkpoints saved every 30 steps. Detailed experimental settings are provided in Appendix A.2. At each checkpoint, we evaluate model performance on a validation set comprising 1,000 questions. For a more detailed analysis, we categorize the validation questions into four distinct difficulty levels: Easy, Medium, Hard, and Hardest. To determine the difficulty level of each question , we prompt the Qwen2.5-Math-7B model with the same question eight times and calculate the average accuracy for that question, denoted as . Based on the value of , each question is assigned to a difficulty level according to the following criteria:
Easy: (i.e., 6, 7, or 8 correct out of 8)
Medium: (i.e., 3, 4, or 5 correct)
Hard: (i.e., 1 or 2 correct)
Hardest: (i.e., 0 correct)
Questions answered correctly at least 6 out of 8 times () are labeled Easy, as the model shows high consistency on these items. Questions never answered correctly () are labeled Hardest, exceeding the model’s capability. Intermediate thresholds ( for Medium and for Hard) are chosen to ensure each category has enough questions for reliable analysis. This classification allows a clear and systematic evaluation of model performance across difficulty levels, with similar numbers of questions in each group.
We monitor the changes in average accuracy during RL and SFT training for Easy, Medium, Hard, and Hardest questions. As shown in Figure 1(a), RL outperforms SFT on Easy and Medium questions, while SFT yields better results than RL on Hardest questions. Notably, after SFT, the LLM fails to solve certain questions that were previously answered correctly by the initial model. In contrast, RL consistently improves accuracy across all difficulty levels and does not exhibit this issue. We also examine changes in the average response length during RL and SFT training, as shown in Figure 1(b). SFT produces longer responses across all question difficulties, aiming to mimic the target long response. In contrast, RL starts with shorter responses, and for medium and hard questions, response length increases over time, indicating more exploration. For the hardest questions, response length remains stable, suggesting the model stops exploring further. Ideally, responses to difficult questions should be longer, while those to simple questions should be shorter. However, SFT and RL each only partially achieve this objective.
We further analyze the evolution of accuracy during training by tracking both the initial and final accuracy for each question, as illustrated in Figure 1(c). Our findings reveal substantial differences in how RL and SFT influence the model’s performance across questions of varying difficulty. For Easy and Medium questions, SFT results in decreased accuracy for some items, whereas RL generally preserves or enhances the model’s original ability to solve these questions. However, this pattern reverses for Hardest questions: SFT enables the model to learn and solve a greater number of these challenging items, as evidenced by the increased transition of Hardest questions into other categories. In contrast, RL does not yield comparable improvements for Hardest questions, highlighting a limitation in its ability to support learning beyond its current capabilities.
Based on these observations, we conclude that accuracy trajectories across questions of varying difficulty do not increase uniformly during RL and SFT training. Instead, a critical distinction emerges between these two paradigms: RL is more effective for Easy and Medium questions, whereas SFT proves more advantageous for the Hardest questions. These findings underscore the complementary strengths of RL and SFT: RL excels at maintaining and improving performance on questions within the model’s original capabilities, while SFT is more effective at facilitating progress on questions that extend beyond those capabilities. Consequently, by selectively incorporating the hardest questions during training and interleaving fine-tuning steps for these questions, it is possible to enhance the model’s overall performance and achieve more scalable and effective RL for LLMs.
2.2 Reinforcement Learning Interleaved with Online Fine-Tuning
Overview. Motivated by the complementary strengths of RL and SFT in addressing questions of varying difficulty, we propose a novel training approach termed Reinforcement Learning Interleaved with Online Fine-Tuning (ReLIFT). ReLIFT dynamically integrates RL and SFT based on the characteristics of the training data and the current capabilities of the model, as illustrated in Figure 2. During RL training, we identify particularly challenging examples based on rollout accuracy. For each of these hardest questions, we obtain high-quality CoT solutions either by collecting them in advance before training or by consulting a stronger model or human annotators, and filter out incorrect answers. These curated examples are added to an SFT buffer. When the buffer accumulates enough challenging questions to form a training batch, we perform an SFT step using these examples. The underlying intuition is that questions difficult for the model to learn through RL alone may benefit from targeted SFT. The timing of SFT is adaptive: when the model performs poorly on training questions, SFT is applied more frequently; as the model improves, RL is emphasized to further enhance its reasoning ability. Below, we provide a detailed description of our method.
Reinforment Learning while Collecting Hard Questions. Owing to the success of Deepseek-R1, GRPO has become the de facto approach for RLVR training. GRPO leverages the reward scores of sampled solutions from a query to estimate the advantage, thereby eliminating the need for an additional value model. Formally, we denote the policy model before and after the update as and . Given a question , a set of solutions generated by , and the reward function , the GRPO objective is defined as follows:
During the rollout phase of GRPO, for each question , we generate a set of solutions , which are evaluated by the reward function to obtain the accuracy . We identify particularly challenging questions as those for which . For these particularly hard questions, we solicit high-quality chain-of-thought (CoT) solutions from a stronger reasoning model (e.g., R1) or human experts, filter out incorrect CoTs to obtain reliable QA pairs , and add them to a buffer . This process is performed online during training and can be formalized as:
where denotes the CoT response obtained either in advance or online from an external model or human annotator, indicates the final answer extracted from , and is the groundtruth answer for . We only keep pairs where .
Interleaved Online Fine-Tuning for Hardest Questions. When the size of the fine-tuning buffer exceeds a predefined threshold (i.e., ), we sample a batch of pairs from the buffer and perform a fine-tuning step using these hard examples. The fine-tuning step minimizes the standard cross-entropy loss:
Notably, the timing of interleaved fine-tuning is adaptive: at the early stage of training, when the model performs poorly on training questions, SFT is applied more frequently to help the model quickly learn better reasoning patterns. As training progresses and the model improves, RL is used more often to further incentivize and enhance the model’s existing reasoning abilities. Moreover, in most cases, it is unnecessary to prepare a large amount of CoTs in advance; instead, it is possible to only generate CoTs online for the hardest questions during training.
3 Experimens
3.1 Experimental Setting
Dataset Construction Our training set is a subset of OpenR1-Math-220k Face (2025), with prompts sourced from NuminaMath 1.5 Li et al. (2024) and detailed demonstrations generated by Deepseek-R1 Guo et al. (2025). We utilize the version of the dataset filtered by LUFFY Yan et al. (2025), resulting in 45,000 prompts and their corresponding CoT demonstrations.
RL and SFT Implementation. For RL, we omit the KL divergence loss term and set the entropy loss coefficient to 0.001. Consistent with Dr.GRPO Liu et al. (2025) and LUFFY Yan et al. (2025), we exclude both length normalization and standard error normalization from the GRPO loss 1 in all experiments. The rollout batch size is set to 128, the rollout times is 8, the update batch size is 64, and the learning rate is . A temperature of 1.0 is used during rollout generation. For SFT, we use a training batch size of 128 and a learning rate of .
ReLIFT Implementation. The RL and fine-tuning configurations for ReLIFT are the same as those described above. During interleaved fine-tuning, we utilize the demonstrations collected in advance. For RL and interleaved SFT, the learning rates are both set to .
Training. By default, we use Qwen2.5-Math-7B Yang et al. (2024), following previous works Yan et al. (2025); Cui et al. (2025); Liu et al. (2025); Zeng et al. ; Hu et al. (2025). To encourage greater exploration, and in line with LUFFY, we increase the RoPE theta from 10,000 to 40,000 and expand the window size to 16,384. We train for 500 steps for both RL and ReLIFT, and train 3 epochs for SFT. Further details are provided in Appendix A.1.
Baseline. (1) Simple-RL Zeng et al. : Trains Qwen2.5-Math-7B using a rule-based reward function. (2) Oat-Zero Liu et al. (2025): Also trains Qwen2.5-Math-7B with rule-based rewards, and further proposes removing the standard deviation in GRPO advantage calculation as well as omitting token-level normalization in the policy loss. (3) PRIME-Zero Cui et al. (2025): Utilizes policy rollouts and outcome labels, leveraging implicit process rewards. (4) OpenReasonerZero Hu et al. (2025): Trains Qwen-2.5-7B with rule-based rewards. (5) LUFFY Yan et al. (2025): Incorporates off-policy reasoning traces into RL by combining them with on-policy rollouts.
Evaluation. For our evaluation, we primarily consider five widely used mathematical reasoning benchmarks: AIME 2024, AIME 2025, AMC Li et al. (2024), OlympiadBench He et al. (2024), and MATH500 Hendrycks et al. (2021). Due to the relatively small test sets for AIME 2024, AIME 2025, and AMC, we report avg@32 for these datasets. For OlympiadBench and MATH500, we use pass@1 as the evaluation metric. Since our reinforcement learning training is mainly centered on mathematical reasoning, we also assess generalization performance on MMLU-Pro Wang et al. (2024). To prevent data contamination, we shuffle the multiple-choice options.
Model | AIME-24 | AIME-25 | AMC | MATH-500 | Olympiad | MMLU-Pro | Overall | |||||||
ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | |
Qwen-Math | 5.1 | 3146 | 1.4 | 3518 | 22.6 | 2611 | 37.2 | 1811 | 19.0 | 2429 | 30.5 | 938 | 19.3 | 2408 |
Qwen-Math-Instruct | 9.7 | 16453 | 8.2 | 15715 | 39.9 | 11734 | 77.2 | 7834 | 34.1 | 12424 | 33.0 | 20919 | 33.7 | 14179 |
SimpleRL-Zero | 27.0 | - | 6.8 | - | 54.9 | - | 76.0 | - | 34.7 | - | 34.5 | - | 38.98 | - |
OpenReasoner-Zero | 16.5 | - | 15.0 | - | 52.1 | - | 82.4 | - | 47.1 | - | 58.7 | - | 45.3 | - |
PRIME-Zero | 17.0 | - | 12.8 | - | 54.0 | - | 81.4 | - | 40.3 | - | 32.7 | - | 39.7 | - |
Oat-Zero | 33.4 | - | 11.9 | - | 61.2 | - | 78.0 | - | 43.4 | - | 41.7 | - | 45.9 | - |
LUFFY | 27.2 | 3120 | 18.1 | 1852 | 61.1 | 1581 | 85.6 | 1370 | 53.6 | 1597 | 52.6 | 1052 | 49.7 | 1762 |
RL | 22.1 | 1131 | 13.9 | 1527 | 59.5 | 449 | 84.2 | 120 | 44.7 | 433 | 48.3 | 260 | 45.5 | 653 |
SFT | 26.9 | 16199 | 23.1 | 16609 | 58.9 | 9397 | 85.0 | 3717 | 50.7 | 9769 | 48.6 | 4906 | 48.9 | 10166 |
ReLIFT | 28.2 | 1666 | 20.1 | 1093 | 64.9 | 943 | 85.2 | 405 | 55.9 | 796 | 52.5 | 1463 | 51.1 | 1061 |
3.2 Main Results
Reasoning Performance. Our main results are presented in Table 1, where we compare ReLIFT with several representative zero-RL baselines, as well as direct SFT, on-policy RL, and off-policy RL method LUFFY Yan et al. (2025). Across five challenging competition-level reasoning benchmarks and one out-of-distribution benchmark, ReLIFT achieves the highest overall accuracy of 51.1%, outperforming all existing baselines and establishing a new state-of-the-art for models based on Qwen2.5-Math-7B. Specifically, ReLIFT exceeds the strongest zero RL baseline (Oat-Zero) by a substantial margin of +5.2 points in overall accuracy, and consistently achieves the best or second-best results on all individual benchmarks. Furthermore, compared to SFT and RL trained on the same dataset, ReLIFT demonstrates clear superiority: it outperforms SFT by +2.2 points and RL by +5.6 points in overall accuracy. ReLIFT also outperforms LUFFY, which combines on-policy and off-policy reasoning traces, despite using only 13% off-policy guidance. These results underscore the effectiveness of integrating RL with interleaved fine-tuning for hardest questions, enabling ReLIFT to achieve both higher performance and better generalization.
Reasoning Efficiency. In addition to accuracy, we evaluate the efficiency of different methods by measuring the average response length (LEN) across all benchmarks. ReLIFT demonstrates remarkable efficiency, generating shorter solutions (average LEN: 1,061) compared to SFT (average LEN: 10,166), Qwen2.5-Math-Instruct (average LEN: 14,179), and LUFFY, while still achieving superior accuracy. This indicates that ReLIFT not only produces more accurate answers but also does so with significantly less computational overhead and more concise reasoning chains.
3.3 Training Dynamics
Training Dynamics of ReLIFT. Figure 3 presents the training dynamics of rewards, response lengths, and the number of hardest questions during RL and ReLIFT training. As shown in Figure 3(a), both the RL and ReLIFT models exhibit an increasing trend in rewards, indicating progressive improvement in performance. Notably, the reward achieved by ReLIFT consistently surpasses that of standard RL, suggesting that ReLIFT is more effective in optimizing the target objective. Regarding response lengths, as depicted in Figure 3(b), RL tends to generate shorter responses as training progresses, which may reflect a diminishing capacity to address more challenging questions. In contrast, ReLIFT demonstrates a gradual increase in response length, implying greater computational effort and a higher potential to solve difficult problems. Furthermore, when examining the number of hardest questions encountered during training (Figure 3(c)), ReLIFT consistently faces fewer such instances compared to RL, and this number decreases over time.
Maintaining Exploration. As shown in Figure 3(d), the entropy of the RL model steadily declines throughout training, indicating reduced exploration. In contrast, the entropy of ReLIFT remains relatively high and exhibits persistent fluctuations, consistently surpassing that of RL. This sustained exploration allows ReLIFT to continue discovering novel solutions, thereby facilitating ongoing performance improvements even in the later stages of training. Actually, our method continues to exhibit a upward trend in reward over the final 500 training steps, underscoring the benefits of sustained exploration for long-term optimization.
3.4 Ablation Study and Analysis
Model | AIME-24 | AIME-25 | AMC | MATH-500 | Olympiad | MMLU-Pro | Overall | |||||||
ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | |
ReLIFT | 28.2 | 1666 | 20.1 | 1093 | 64.9 | 943 | 85.2 | 405 | 55.9 | 796 | 52.5 | 1463 | 51.1 | 1061 |
ReLIFT(all) | 7.8 | 26583 | 9.0 | 26704 | 31.6 | 22025 | 54.2 | 16984 | 26.4 | 21670 | 13.7 | 27938 | 23.8 | 23650 |
ReLIFT(uniform) | 23.2 | 3103 | 19.3 | 2762 | 62.3 | 1389 | 86.0 | 962 | 52.4 | 1426 | 51.4 | 655 | 49.1 | 1716 |
ReLIFT(random) | 22.1 | 1131 | 13.9 | 1527 | 59.5 | 449 | 84.2 | 120 | 44.7 | 433 | 48.4 | 22644 | 45.5 | 4367 |
We conduct an ablation study to assess the effectiveness of our ReLIFT approach. Specifically, we compare the following three ablation settings:
ReLIFT(all): Alternates one fine-tuning step after each RL step on the same batch.
ReLIFT(uniform): Applies fine-tuning at fixed, uniform intervals (every 8 RL steps), matching the number of demonstrations with ReLIFT.
ReLIFT(random): Performs interleaved FT. The buffer is populated based on the number of Hardest questions, but instead of using the hardest questions, it is randomly filled with other, non-hardest questions.
As summarized in Table 2, the results show that ReLIFT (all) quickly collapses, probably due to conflicting optimization objectives between RL and FT. ReLIFT (uniform) not only reduces accuracy but also leads to longer responses compared to ReLIFT. ReLIFT (random) causes a dramatic increase in average response length during training, while still underperforming ReLIFT. We also analyze the number of demonstrations of different methods and SFT, as depicted in Figure 4. ReLIFT requires only 13% of the demonstrations used in SFT.
Overall, these findings demonstrate that simply alternating RL and fine-tuning is ineffective; both the scheduling and the selection of fine-tuning training data are critical. ReLIFT applies fine-tuning more frequently in the early stages, when the model’s reasoning ability is relatively weak, and focuses on the hardest questions. This enables the model to efficiently acquire new knowledge without sacrificing performance or conciseness. In contrast, all, uniform, or random fine-tuning scheduling leads to suboptimal results and unnecessary inflation of response length.
3.5 Extension to More Models
Model | AIME-24 | AIME-25 | AMC | MATH-500 | Olympiad | MMLU-Pro | Overall | |||||||
ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | |
Qwen-Math-1.5B | 3.0 | 4105 | 1.4 | 4291 | 19.4 | 3480 | 44.4 | 2637 | 16.7 | 3487 | 5.0 | 6260 | 15.0 | 4043 |
Qwen-Math-1.5B-Instruct | 8.1 | 5722 | 6.6 | 4468 | 36.9 | 4614 | 66.0 | 6148 | 30.7 | 5526 | 24.2 | 8360 | 28.8 | 5806 |
SFT | 12.7 | 18060 | 13.0 | 19271 | 40.9 | 12538 | 71.8 | 5879 | 33.8 | 12906 | 23.5 | 11767 | 32.6 | 13403 |
RL | 10.2 | 1487 | 8.7 | 1187 | 44.7 | 796 | 71.4 | 423 | 34.5 | 832 | 28.7 | 756 | 33.0 | 913 |
ReLIFT | 13.1 | 6294 | 8.8 | 5866 | 42.5 | 3856 | 73.6 | 2750 | 36.0 | 3937 | 30.1 | 3969 | 34.0 | 4421 |
Qwen-7B | 5.1 | 3146 | 1.4 | 3518 | 22.6 | 2611 | 37.2 | 1811 | 19.0 | 2429 | 39.8 | 434 | 20.9 | 2324 |
Qwen-7B-Instruct | 10.3 | 1299 | 6.5 | 562 | 34.9 | 1043 | 67.6 | 605 | 29.8 | 907 | 53.3 | 7413 | 33.7 | 1804 |
SFT | 15.7 | 19363 | 18.6 | 18070 | 49.8 | 11349 | 80.8 | 4453 | 44.3 | 11528 | 54.6 | 5188 | 43.9 | 11658 |
RL | 13.3 | 4371 | 4.2 | 3988 | 40.2 | 2845 | 65.4 | 1971 | 31.4 | 2852 | 56.6 | 180 | 35.2 | 2867 |
ReLIFT | 19.1 | 12197 | 14.7 | 9284 | 51.9 | 5909 | 81.6 | 2395 | 46.1 | 6160 | 56.4 | 2550 | 45.0 | 6415 |
To evaluate the general applicability of our method, we extend ReLIFT to both a smaller base model, Qwen2.5-Math-1.5B Yang et al. (2024), and a weaker base model, Qwen2.5-7B Team (2024). Table 3 presents the performance across five challenging competition-level math benchmarks and one out-of-domain benchmark. On both models, ReLIFT consistently achieves substantial improvements, surpassing both SFT and RL. Across all benchmarks, our method ranks first or second, while maintaining an acceptable response length. These results underscore the applicability of ReLIFT to achieve both higher performance and better generalization.
4 Related Works
4.1 RL-Driven Emergence of Reasoning Abilities
Reinforcement learning has recently emerged as a promising approach for enhancing the reasoning capabilities of large language models (LLMs). Early breakthroughs, such as O1 Jaech et al. (2024), DeepSeek R1 Guo et al. (2025), and Kimi 1.5 Team et al. (2025), have demonstrated that applying Reinforcement Learning with Verifiable Rewards (RLVR) can lead to significant improvements in complex reasoning tasks. Notably, DeepSeek R1 further shows that applying RLVR directly to a base model—without the need for intermediate supervised fine-tuning—can also effectively enhance the reasoning abilities of LLMs.
Building on these foundations, Light-r1 Xie et al. (2025) incorporated curriculum learning to further refine reasoning skills, while Logic-rl Xie et al. (2025) adopted logic-driven reward functions to improve general reasoning ability. Deepscaler Luo et al. (2025) introduced Iterative Context Lengthening, a method that trains models with progressively longer contexts as their performance improves, thereby reducing both monetary costs and end-to-end training time. In addition, Oat-Zero Liu et al. (2025) was trained from Qwen2.5-Math-7B using rule-based rewards, and proposed removing the standard deviation in GRPO advantage computation as well as token-level normalization in policy loss computation. PRIME-Zero Cui et al. (2025) leveraged policy rollouts and outcome labels, utilizing implicit process rewards to further enhance model performance.
4.2 Training Paradigms
Despite significant advances in reinforcement learning (RL), Yue et al. Yue et al. (2025) argue that RL with value-based rewards (RLVR) does not enable large language models (LLMs) to acquire genuinely new reasoning abilities beyond those present in their base models. Instead, RL primarily enhances performance by biasing the model toward reasoning paths it already knows that are more likely to yield rewards. In other words, most RL methods optimize within the model’s existing knowledge rather than facilitating the acquisition of new information, thereby limiting the LLM’s overall capabilities.
In contrast, directly fine-tuning on high-quality demonstration data can introduce new knowledge and abilities to the model, and this approach is widely used to adapt LLMs to downstream reasoning tasks. However, supervised fine-tuning (SFT) requires high-quality, detailed demonstration data and tends to encourage memorization of the training data, often resulting in poor generalization to out-of-distribution scenarios Chu et al. (2025).
Some approaches attempt to combine supervised fine-tuning (SFT) and reinforcement learning (RL). Typically, these methods first use SFT to stabilize the model’s output format, then apply RL to further improve performance and potentially unlock more advanced reasoning abilities in LLMs Guo et al. (2025); Wen et al. (2025). However, such combined methods remain limited: SFT is highly dependent on the quality of demonstration data, and the subsequent RL phase is still confined to the knowledge already present in the SFT-trained model, making it difficult for the model to acquire genuinely new information.
LUFFY Yan et al. (2025) seeks to address these limitations by augmenting RL with off-policy reasoning traces, dynamically balancing imitation and exploration through a combination of off-policy demonstrations and on-policy rollouts during training. However, this approach requires the preparation of off-policy data in advance and relies heavily on probability-sharpening techniques, which may restrict its applicability. In contrast, our work proposes a simpler and more direct method that combines RL and SFT to leverage the strengths of both paradigms. Our approach encourages LLMs to explore when possible and to incorporate new information as needed, thereby pushing RL-driven reasoning beyond the boundaries of existing knowledge.
5 Conclusion
In summary, we systematically analyze the respective strengths of RL and SFT for LLM reasoning, finding that RL is more effective for learning easier questions while SFT is crucial for solving challenging questions. Building on these insights, we introduce ReLIFT, an approach that interleaves RL with online Fine-Tuning on the hardest questions, achieving state-of-the-art results with significantly less demonstration data and more concise solutions. Future work will focus on scaling ReLIFT to larger models and developing more effective strategies for coordinating SFT and RL to further advance LLM reasoning and generalization.
Appendix A Experiment Details
A.1 Main Experiment Details
Training. In addition to Qwen2.5-Math-7B, we extend ReLIFT to both Qwen2.5-Math-1.5B Yang et al. (2024) and Qwen2.5-7B Team (2024). Since the maximum sequence length for Qwen2.5-Math-1.5B is 4096, which is insufficient for our tasks, we follow the approach used for Qwen2.5-Math-7B by increasing the RoPE theta from 10,000 to 40,000 and expanding the window size to 16,384. To ensure fairness, we maintain 8 rollouts per prompt for all RL-trained models. The learning rate is fixed at . All training experiments are conducted on 2 8 A800 GPUs. We train all RL models for 500 steps and all SFT models for three epochs. For all experiments, we use verl Sheng et al. (2024).
Evaluation. All evaluations are conducted using vllm Kwon et al. (2023), with the temperature set to 0.6 and the maximum number of tokens set to 8192. Results are verified using math-verify Face (2024). For the baselines, we independently reproduce and verify the results of the LUFFY Yan et al. (2025), whereas the results for the other zero-RL models are reported as in the LUFFY paper.
Chat Template. To encourage the base model to reason more effectively, following Yan et al. (2025), we adopt a complex chat template as shown in Figure 5. All training experiments and evaluation utilize the same chat template.
Reward Function. To evaluate the impact of our method, we adopt a simple reward function as below. All training experiments employ the same reward function.
A.2 Motivation Experiment Details
Dataset. We randomly sample a subset of 8,000 examples from the training dataset described in 3.1 as the training set, and select 1,000 examples as the validation set. RL and SFT are both conducted on this subset dataset for 120 steps, using the same parameters as in our main experiments.

### v2

@@
 Abstract
-Recent advances in large language model (LLM) reasoning have shown that sophisticated behaviors such as planning and self-reflection can emerge through reinforcement learning (RL).
-However, despite these successes, RL in its current form remains insufficient to induce capabilities that exceed the limitations of the base model, as it is primarily optimized based on existing knowledge of the model rather than facilitating the acquisition of new information.
+Recent advances in large language model (LLM) reasoning have shown that reasoning ability can emerge through reinforcement learning (RL).
+However, despite these successes, RL in its current form remains insufficient to induce capabilities that exceed the limitations of the base model, as it is primarily optimized based on existing knowledge of the model.
 To address this limitation, we employ supervised fine-tuning (SFT) to learn what RL cannot, which enables the incorporation of new knowledge and reasoning patterns by leveraging high-quality demonstration data.
-We analyze the training dynamics of RL and SFT for LLM reasoning and find that RL excels at maintaining and improving performance on questions within the model’s original capabilities, while SFT is more effective at enabling progress on questions beyond the current scope of the model.
-Motivated by the complementary strengths of RL and SFT, we introduce a novel training approach, ReLIFT (Reinforcement Learning Interleaved with Online Fine-Tuning).
-In ReLIFT, the model is primarily trained using RL, but when it encounters challenging questions, high-quality solutions are collected for fine-tuning, and the training process alternates between RL and fine-tuning to enhance the model’s reasoning abilities.
-ReLIFT achieves an average improvement of over +5.2 points across five competition-level benchmarks and one out-of-distribution benchmark compared to other zero-RL models.
-Furthermore, we demonstrate that ReLIFT outperforms both RL and SFT while using only 13% of the detailed demonstration data, highlighting its scalability.
-These results provide compelling evidence that ReLIFT overcomes the fundamental limitations of RL and underscores the significant potential.
-Project Page: https://github.com/TheRoadQaQ/ReLIFT
+We analyze the training dynamics of RL and SFT for LLM reasoning and find that RL excels at improving performance on questions within the model’s original capabilities, while SFT is more effective at enabling progress on questions beyond the current scope of the model.
+Motivated by the complementary strengths of RL and SFT, we introduce ReLIFT (Reinforcement Learning Interleaved with Online Fine-Tuning), a novel training strategy.
+ReLIFT employs RL for general training, but interleaves it with targeted SFT on challenging questions for which high-quality solutions are collected online.
+By alternating between RL and SFT, ReLIFT addresses model weaknesses as they emerge.
+Empirically, ReLIFT outperforms previous RLVR methods by an average of +6.7 points across a suite of six benchmarks (five math reasoning and one out-of-distribution).
+More importantly, ReLIFT surpasses baselines such as individual RL, individual SFT, and various hybrid approaches while reducing the required training time.
+These results provide compelling evidence that ReLIFT is a powerful and resource-efficient paradigm for developing capable reasoning models.
+The code is available at here.
 1 Introduction
-Recent advancements in large language model (LLM) reasoning, as demonstrated by OpenAI-o1 Jaech et al.
+Recent advancements in large language model reasoning, as demonstrated by OpenAI o-series Jaech et al.
 (2024), DeepSeek-R1 Guo et al.
-(2025), and Kimi-1.5 Team et al.
-(2025), highlight the potential to enhance reasoning capabilities by increasing test-time computational resources.
+(2025), Gemini 2.5 Comanici et al.
+(2025), and Kimi k-series Team et al.
+(2025b), highlight the potential to enhance reasoning capabilities by increasing test-time computational resources.
 These models typically generate long Chains-of-Thought (CoT Wei et al.
 (2022)) responses, exhibiting sophisticated behaviors such as reflection Kumar et al.
-(2024); Team et al.
-(2025) and planning Team et al.
+(2024) and planning Ke et al.
 (2025).
-The primary factor driving this progress is large-scale Reinforcement Learning with Verifiable Rewards (RLVR), which does not depend on traditional scaffolding techniques like Monte Carlo Tree Search Xie et al.
+The primary factor driving this progress is large-scale Reinforcement Learning with Verifiable Rewards (RLVR), which does not depend on traditional techniques like Monte Carlo Tree Search Xie et al.
 (2024); Li et al.
 (2025) or Process Reward Models Setlur et al.
+(2024); Zhang et al.
 (2025).
-In RLVR, rewards are assigned based on whether the model’s output matches a ground-truth solution in mathematics or passes unit tests in code, thereby enabling scalability without the need for detailed demonstration data.
-In practice, RLVR is typically implemented using policy optimization algorithms such as Proximal Policy Optimization (PPO) Schulman et al.
+In RLVR, rewards are assigned based on whether the model’s output matches a ground-truth solution in mathematics or passes unit tests in code, thereby enabling scalability without the need for demonstration data.
+Implemented with policy optimization algorithms like Proximal Policy Optimization (PPO) Schulman et al.
 (2017) and Group Relative Policy Optimization (GRPO) Shao et al.
-(2024).
-Large-scale RLVR, characterized by its simplicity and directness, has enabled the emergence of long CoT reasoning and self-reflection capabilities.
-However, despite the empirical successes of RLVR, its current form is insufficient to incentivize capabilities that transcend the base model’s limitations Yue et al.
+(2024), the simplicity and directness of large-scale RLVR have been pivotal in incentivizing the reasoning capabilities seen in today’s most advanced LLMs.
+Despite its empirical successes, RLVR is insufficient to foster capabilities that transcend a base model’s inherent limitations Yue et al.
+(2025); Cheng et al.
 (2025); Zhao et al.
 (2025).
-RL is inherently on-policy, learning from the LLM’s self-generated responses through iterative rollouts and feedbacks.
-Essentially, RL primarily enhances performance by biasing the model toward reasoning paths it already knows are more likely to yield rewards.
-In other words, most RL methods optimize based on the model’s existing knowledge rather than facilitating the acquisition of new information.
+Research suggests that RLVR primarily reinforces existing behaviors rather than instilling novel reasoning abilities.
+For instance, Yue et al.
+(2025) argue that RLVR fails to equip LLMs with new reasoning skills, noting that RLVR-trained models underperform base models on pass@k at higher k values, indicating limited expansion of reasoning capabilities.
+Similarly, Zhao et al.
+(2025) show that RL post-training amplifies certain pre-training behaviors while suppressing others.
+Furthermore, Cheng et al.
+(2025) find that RL stifles exploration, causing models to converge on narrow behavioral patterns and plateau on complex reasoning tasks.
+This limitation is rooted in the on-policy nature of RL, which learns from the model’s own generated responses via iterative rollouts and feedback.
+Consequently, RL primarily enhances performance by biasing the model toward reasoning paths it already knows are likely to yield rewards, optimizing existing knowledge rather than facilitating the acquisition of new information.
 In contrast, supervised fine-tuning (SFT) offers a complementary approach to enhancing the reasoning capabilities of LLMs.
@@
 (2024); Yue et al.
-(2025), and has been shown to achieve superior reasoning performance compared to RL, particularly in small-scale models Guo et al.
+(2025); Guo et al.
+(2025), and has been shown to achieve superior reasoning performance compared to RL, particularly in small-scale LLMs Guo et al.
 (2025).
-Nevertheless, the effectiveness of SFT is highly contingent upon the availability of substantial amounts of high-quality demonstration data.
-Furthermore, models trained with SFT often exhibit limited generalization to out-of-distribution (OOD) scenarios, whereas RL-based approaches tend to demonstrate greater robustness Chu et al.
+Nevertheless, the effectiveness of SFT relies highly on the availability of substantial amounts of high-quality demonstration data.
+Furthermore, SFT-trained models often exhibit limited generalization to out-of-distribution (OOD) scenarios, while RL-based approaches tend to demonstrate greater robustness Chu et al.
+(2025); Chen et al.
 (2025).
@@
 How can RL and SFT be effectively combined to improve LLM reasoning and OOD generalization, reduce dependence on expensive demonstration data, and achieve capabilities beyond current cognitive constraints?
-To achieve this goal, we analyze the training dynamics by examining how the accuracy of the questions change under both RL and SFT.
-In Section 2.1, we evaluate the accuracy of the model on the questions at various checkpoints and classify these questions into four levels of difficulty.
-Our findings show that RL is more effective for relatively easy questions, whereas SFT is more beneficial for most difficult ones.
-Furthermore, for simple questions, SFT can degrade the model’s existing performance, as accuracy on these questions decreases and response length increases.
+To achieve this goal, we analyze the training dynamics of RL and SFT by examining how the accuracy of questions changes throughout the training process.
+We evaluate the model’s accuracy on questions at various checkpoints, and classify them into four levels of difficulty.
+Our findings indicate that RL is more effective for questions of lower difficulty, whereas SFT proves more beneficial for the most challenging ones.
+Furthermore, for simpler questions, SFT can degrade the model’s existing performance and the response length increases.
 In contrast, for challenging questions, the improvement provided by RL is less pronounced compared to SFT.
@@
 Once the number of challenging questions in the SFT buffer is sufficient, we perform SFT on these questions for one step.
-We conduct comprehensive experiments on five competition-level math reasoning benchmarks and one out-of-distribution (OOD) benchmark.
-Our proposed ReLIFT method achieves a new state-of-the-art accuracy of 51.1% based on Qwen2.5-Math-7B, outperforming previous zero-RL baselines by a significant margin, surpassing the strongest zero RL baseline by +5.2 points.
-Notably, ReLIFT demonstrates clear advantages over both pure SFT and pure RL, while requiring only 13% amount of detailed demonstration data.
-In addition, ReLIFT generates much more concise solutions, reducing average response length by nearly 10x compared to SFT, thus improving both performance and efficiency.
-We further validate the generality of ReLIFT by extending it to smaller and weaker foundation models, consistently observing superior results over standard RL and SFT approaches.
+We conduct comprehensive experiments on five challenging math reasoning benchmarks and one OOD benchmark.
+Our proposed ReLIFT method achieves a new state-of-the-art accuracy of 52.6% with Qwen2.5-Math-7B, which significantly outperforms previous RLVR baselines.
+Notably, ReLIFT demonstrates clear superiority over methods including pure SFT, pure RL, and combined RL and SFT approaches, such as RL with SFT loss, SFT followed by RL, and LUFFY Yan et al.
+(2025).
+ReLIFT requires minimal detailed demonstration data and GPU training hours.
+Additionally, ReLIFT generates significantly more concise solutions, which improves both performance and efficiency.
+We further validate the generality of ReLIFT by extending it to smaller and weaker base models, consistently observing superior results.
+The main contributions of this work are as follows:
+We present a systematic analysis of the training dynamics of RL and SFT in reasoning, empirically demonstrating their complementary roles: RL refines existing skills on solvable problems, while SFT is essential for acquiring new knowledge for difficult ones.
+We propose ReLIFT, a novel framework that interleaves RL with online fine-tuning.
+ReLIFT dynamically identifies challenging examples during RL for targeted fine-tuning to address emerging model weaknesses.
+Extensive experiments show ReLIFT achieves state-of-the-art performance on mathematical reasoning and OOD benchmarks, outperforming pure SFT, pure RL, and existing hybrid methods with significantly less demonstration data and computational overhead.
 2 Reinforcement Learning Interleaved with Online Fine-Tuning
 2.1 RL vs. SFT: Training Dynamics Across Question Difficulty Levels
-We start our study by examining how the model’s accuracy evolves on questions of varying difficulty during both SFT and RL training.
-Specifically, we independently perform SFT and RL on the Qwen2.5-Math-7B model, using a subset of the Open-R1-math-220K Face (2025).
+We begin our study by examining how the model’s accuracy evolves on questions of varying difficulty during both SFT and RL training.
+Specifically, we independently train the Qwen2.5-Math-7B model using both SFT and RL on a subset of the Open-R1-math-220K dataset Face (2025).
 Both SFT and RL are conducted for 120 steps, with model checkpoints saved every 30 steps.
-Detailed experimental settings are provided in Appendix A.2.
+Detailed experimental settings are provided in Appendix B.
 At each checkpoint, we evaluate model performance on a validation set comprising 1,000 questions.
-For a more detailed analysis, we categorize the validation questions into four distinct difficulty levels: Easy, Medium, Hard, and Hardest.
-To determine the difficulty level of each question , we prompt the Qwen2.5-Math-7B model with the same question eight times and calculate the average accuracy for that question, denoted as .
-Based on the value of , each question is assigned to a difficulty level according to the following criteria:
-Easy: (i.e., 6, 7, or 8 correct out of 8)
-Medium: (i.e., 3, 4, or 5 correct)
-Hard: (i.e., 1 or 2 correct)
-Hardest: (i.e., 0 correct)
-Questions answered correctly at least 6 out of 8 times () are labeled Easy, as the model shows high consistency on these items.
-Questions never answered correctly () are labeled Hardest, exceeding the model’s capability.
-Intermediate thresholds ( for Medium and for Hard) are chosen to ensure each category has enough questions for reliable analysis.
-This classification allows a clear and systematic evaluation of model performance across difficulty levels, with similar numbers of questions in each group.
-We monitor the changes in average accuracy during RL and SFT training for Easy, Medium, Hard, and Hardest questions.
-As shown in Figure 1(a), RL outperforms SFT on Easy and Medium questions, while SFT yields better results than RL on Hardest questions.
-Notably, after SFT, the LLM fails to solve certain questions that were previously answered correctly by the initial model.
+To facilitate a more detailed analysis, we categorize the validation questions into four distinct difficulty levels: Easy, Medium, Hard, and Hardest.
+To determine the difficulty level of each question , we prompt the Qwen2.5-Math-7B model eight times and calculate the average accuracy for that question, denoted as .
+Following the approach in works like Wen et al.
+(2025); Bae et al.
+(2025); Yu et al.
+(2025), and to ensure a sufficient number of questions in each category, we select these specific thresholds to enable a robust analysis of the training dynamics of both RL and SFT.
+The criteria for assigning difficulty levels are as follows:
+Easy: 6 to 8 correct answers out of 8 ().
+Medium: 3 to 5 correct answers out of 8 ().
+Hard: 1 or 2 correct answers out of 8 ().
+Hardest: 0 correct answers out of 8 (), representing items that are beyond the model’s current capabilities.
+We analyze the average accuracy of LLM trained with RL and SFT on four difficulty levels of questions: Easy, Medium, Hard, and Hardest.
+As shown in Figure 1(a), RL outperforms SFT on Easy and Medium questions.
+Conversely, SFT achieves better results than RL on the Hardest questions.
+Notably, after SFT, the LLM fails to solve certain questions that the initial model previously answered correctly.
 In contrast, RL consistently improves accuracy across all difficulty levels and does not exhibit this issue.
@@
 SFT produces longer responses across all question difficulties, aiming to mimic the target long response.
-In contrast, RL starts with shorter responses, and for medium and hard questions, response length increases over time, indicating more exploration.
-For the hardest questions, response length remains stable, suggesting the model stops exploring further.
-Ideally, responses to difficult questions should be longer, while those to simple questions should be shorter.
-However, SFT and RL each only partially achieve this objective.
+In contrast, RL begins with shorter responses, and for Medium and Hard questions, the response length increases over time, indicating more exploration.
+For the Hardest questions, the response length remains stable, which suggests the model stops exploring further.
+Ideally, responses to difficult questions should be longer, and responses to simple questions should be shorter.
+However, neither SFT nor RL fully achieves this objective.
 We further analyze the evolution of accuracy during training by tracking both the initial and final accuracy for each question, as illustrated in Figure 1(c).
 Our findings reveal substantial differences in how RL and SFT influence the model’s performance across questions of varying difficulty.
-For Easy and Medium questions, SFT results in decreased accuracy for some items, whereas RL generally preserves or enhances the model’s original ability to solve these questions.
+For Easy and Medium questions, SFT results in decreased accuracy for some items, whereas RL generally preserves or enhances the LLM’s original ability to solve these questions.
 However, this pattern reverses for Hardest questions: SFT enables the model to learn and solve a greater number of these challenging items, as evidenced by the increased transition of Hardest questions into other categories.
 In contrast, RL does not yield comparable improvements for Hardest questions, highlighting a limitation in its ability to support learning beyond its current capabilities.
-Based on these observations, we conclude that accuracy trajectories across questions of varying difficulty do not increase uniformly during RL and SFT training.
-Instead, a critical distinction emerges between these two paradigms: RL is more effective for Easy and Medium questions, whereas SFT proves more advantageous for the Hardest questions.
-These findings underscore the complementary strengths of RL and SFT: RL excels at maintaining and improving performance on questions within the model’s original capabilities, while SFT is more effective at facilitating progress on questions that extend beyond those capabilities.
-Consequently, by selectively incorporating the hardest questions during training and interleaving fine-tuning steps for these questions, it is possible to enhance the model’s overall performance and achieve more scalable and effective RL for LLMs.
+These findings demonstrate that RL and SFT have complementary strengths.
+RL excels at refining and preserving the model’s existing knowledge, making it suitable for questions within its competence.
+However, SFT is superior for acquiring new knowledge to tackle problems that lie beyond the model’s initial capabilities.
+This suggests that accuracy does not improve uniformly and that a hybrid training strategy could be optimal.
+By selectively interleaving SFT steps for the Hardest questions, it may be possible to enhance overall model performance and develop more scalable and effective RL for LLMs.
 2.2 Reinforcement Learning Interleaved with Online Fine-Tuning
@@
 Motivated by the complementary strengths of RL and SFT in addressing questions of varying difficulty, we propose a novel training approach termed Reinforcement Learning Interleaved with Online Fine-Tuning (ReLIFT).
-ReLIFT dynamically integrates RL and SFT based on the characteristics of the training data and the current capabilities of the model, as illustrated in Figure 2.
-During RL training, we identify particularly challenging examples based on rollout accuracy.
-For each of these hardest questions, we obtain high-quality CoT solutions either by collecting them in advance before training or by consulting a stronger model or human annotators, and filter out incorrect answers.
-These curated examples are added to an SFT buffer.
-When the buffer accumulates enough challenging questions to form a training batch, we perform an SFT step using these examples.
-The underlying intuition is that questions difficult for the model to learn through RL alone may benefit from targeted SFT.
-The timing of SFT is adaptive: when the model performs poorly on training questions, SFT is applied more frequently; as the model improves, RL is emphasized to further enhance its reasoning ability.
+ReLIFT dynamically integrates RL and fine-tuning (FT) based on the characteristics of the training data and the current capabilities of the model, as illustrated in Figure 2.
+During RL training, we identify hardest examples based on rollout accuracy.
+For each of these hardest questions, we obtain high-quality CoT solutions either by collecting them in advance before training or by consulting a stronger model or human annotators.
+These curated examples are added to a fine-tuning buffer.
+When the buffer accumulates enough questions to form a training batch, we perform a fine-tuning step using these examples.
+The underlying intuition is that questions difficult for the model to learn through RL alone may benefit from targeted fine-tuning.
 Below, we provide a detailed description of our method.
@@
 We identify particularly challenging questions as those for which .
-For these particularly hard questions, we solicit high-quality chain-of-thought (CoT) solutions from a stronger reasoning model (e.g., R1) or human experts, filter out incorrect CoTs to obtain reliable QA pairs , and add them to a buffer .
-This process is performed online during training and can be formalized as:
+For these particularly hard questions, we solicit high-quality CoT solutions from a stronger reasoning model (e.g., DeepSeek-R1) or human experts, filter out incorrect CoTs to obtain reliable QA pairs , and add them to a buffer .
+The data collected from is then added to a primary fine-tuning buffer, , which serves as the data for subsequent fine-tuning step.
+This process is performed online during training.
+Collecting and training can be done together, and can be formalized as:
 where denotes the CoT response obtained either in advance or online from an external model or human annotator, indicates the final answer extracted from , and is the groundtruth answer for .
@@
 The fine-tuning step minimizes the standard cross-entropy loss:
-Notably, the timing of interleaved fine-tuning is adaptive: at the early stage of training, when the model performs poorly on training questions, SFT is applied more frequently to help the model quickly learn better reasoning patterns.
-As training progresses and the model improves, RL is used more often to further incentivize and enhance the model’s existing reasoning abilities.
-Moreover, in most cases, it is unnecessary to prepare a large amount of CoTs in advance; instead, it is possible to only generate CoTs online for the hardest questions during training.
-3 Experimens
-3.1 Experimental Setting
-Dataset Construction Our training set is a subset of OpenR1-Math-220k Face (2025), with prompts sourced from NuminaMath 1.5 Li et al.
-(2024) and detailed demonstrations generated by Deepseek-R1 Guo et al.
+To prevent fine-tuning updates from overly constraining the model’s exploratory behavior, we incorporate an entropy regularization term into the loss function. The full objective is as follows:
+where . The term represents the entropy of the token generation distribution , and is a hyperparameter controlling the weight of the entropy term.
+The frequency of this interleaved fine-tuning is adaptive.
+Early in training, when the model’s performance is low, SFT is applied more frequently to accelerate the learning of effective reasoning patterns.
+As training progresses and the model improves, RL is prioritized to further incentivize the model’s existing abilities.
+Furthermore, this online approach obviates the need for a large, pre-collected dataset of CoT demonstrations .
+Instead, CoTs can be generated dynamically only for the most challenging questions encountered during training.
+3 Experimental Setup
+Dataset Construction.
+Our training set is a subset of OpenR1-Math-220k Face (2025), with prompts sourced from NuminaMath 1.5 Li et al.
+(2024) and detailed demonstration data generated by Deepseek-R1 Guo et al.
 (2025).
-We utilize the version of the dataset filtered by LUFFY Yan et al.
-(2025), resulting in 45,000 prompts and their corresponding CoT demonstrations.
-RL and SFT Implementation.
-For RL, we omit the KL divergence loss term and set the entropy loss coefficient to 0.001.
-Consistent with Dr.GRPO Liu et al.
-(2025) and LUFFY Yan et al.
-(2025), we exclude both length normalization and standard error normalization from the GRPO loss 1 in all experiments.
-The rollout batch size is set to 128, the rollout times is 8, the update batch size is 64, and the learning rate is .
-A temperature of 1.0 is used during rollout generation.
-For SFT, we use a training batch size of 128 and a learning rate of .
-ReLIFT Implementation.
-The RL and fine-tuning configurations for ReLIFT are the same as those described above.
-During interleaved fine-tuning, we utilize the demonstrations collected in advance.
-For RL and interleaved SFT, the learning rates are both set to .
-Training.
-By default, we use Qwen2.5-Math-7B Yang et al.
-(2024), following previous works Yan et al.
-(2025); Cui et al.
-(2025); Liu et al.
-(2025); Zeng et al.
-; Hu et al.
-(2025).
-To encourage greater exploration, and in line with LUFFY, we increase the RoPE theta from 10,000 to 40,000 and expand the window size to 16,384.
-We train for 500 steps for both RL and ReLIFT, and train 3 epochs for SFT.
-Further details are provided in Appendix A.1.
-Baseline.
-(1) Simple-RL Zeng et al.
-: Trains Qwen2.5-Math-7B using a rule-based reward function.
-(2) Oat-Zero Liu et al.
-(2025): Also trains Qwen2.5-Math-7B with rule-based rewards, and further proposes removing the standard deviation in GRPO advantage calculation as well as omitting token-level normalization in the policy loss.
-(3) PRIME-Zero Cui et al.
-(2025): Utilizes policy rollouts and outcome labels, leveraging implicit process rewards.
-(4) OpenReasonerZero Hu et al.
-(2025): Trains Qwen-2.5-7B with rule-based rewards.
-(5) LUFFY Yan et al.
-(2025): Incorporates off-policy reasoning traces into RL by combining them with on-policy rollouts.
+We use the default 94k prompt subset and filter out generations that are longer than 8192 tokens or are verified incorrect by Math-verify Face (2024).
+This process yields 46k prompts and high-quality demonstrations, which we term OpenR1-Math-46k-8192.
 Evaluation.
-For our evaluation, we primarily consider five widely used mathematical reasoning benchmarks: AIME 2024, AIME 2025, AMC Li et al.
+For our evaluation, we primarily consider five widely adopted mathematical reasoning benchmarks: AIME 2024, AIME 2025, AMC Li et al.
 (2024), OlympiadBench He et al.
@@
 To prevent data contamination, we shuffle the multiple-choice options.
+A testing temperature of 0.6 was used for all evaluations.
+Baselines.
+We benchmark ReLIFT against the following baselines using Qwen2.5-Math-7B Yang et al.
+(2024) as the base model.
+For SFT methods, we fine-tuning the base model on the OpenR1-Math-46k-8192 dataset.
+For RL methods, we include GRPO Shao et al.
+(2024) trained on the same 46k dataset; SimpleRL-Zero Zeng et al.
+(2025), which applies GRPO to approximately 24k mathematical samples from GSM8K Cobbe et al.
+(2021) and MATH dataset Hendrycks et al.
+(2021); OpenReasoner-Zero Hu et al.
+(2025), a PPO-based approach trained on 129k multi-source samples; PRIME-Zero Cui et al.
+(2025), which conducts policy rollouts on 150k NuminaMath queries with implicit process rewards; and Oat-Zero Liu et al.
+(2025), which removes the standard deviation in GRPO advantage calculation, and is trained on the MATH dataset.
+For methods combining SFT and RL, we evaluate three approaches: RL w/ SFT loss, which incorporates a SFT loss term into the GRPO objective trained on the 46k dataset; LUFFY Yan et al.
+(2025), a mixed-policy GRPO approach that also utilizes the 46k dataset; and SFT then RL, a two-stage method that applies RL training after an initial SFT phase.
+Specifically, we evaluate two versions of SFT then RL: SFT then RL (v1) involves taking our trained SFT model and training it with RL for 300 steps on a held-out split of the OpenR1-Math-220k dataset, which contains about 49,000 prompts, similar to previous work Cui et al.
+(2025); Yan et al.
+(2025).
+For SFT then RL (v2), to fairly compare with ReLIFT, which uses 8,640 demonstration samples, we first perform SFT on a subset of 8,640 samples from OpenR1-Math-46k-8192.
+Then we conduct GRPO training on the full OpenR1-Math-46k-8192 dataset.
+For a more detailed setup for training these baselines, please refer to Appendix B.
+4 Experimental Results
 | Model | AIME-24 | AIME-25 | AMC | MATH-500 | Olympiad | MMLU-Pro | Overall | |||||||
 | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | |
-Qwen-Math | 5.1 | 3146 | 1.4 | 3518 | 22.6 | 2611 | 37.2 | 1811 | 19.0 | 2429 | 30.5 | 938 | 19.3 | 2408 |
-Qwen-Math-Instruct | 9.7 | 16453 | 8.2 | 15715 | 39.9 | 11734 | 77.2 | 7834 | 34.1 | 12424 | 33.0 | 20919 | 33.7 | 14179 |
-SimpleRL-Zero | 27.0 | - | 6.8 | - | 54.9 | - | 76.0 | - | 34.7 | - | 34.5 | - | 38.98 | - |
+| Qwen2.5-Math-7B | ||||||||||||||
+| Qwen-Math | 16.4 | 2125 | 6.9 | 2039 | 45.5 | 1493 | 65.2 | 1171 | 30.1 | 1594 | 24.9 | 830 | 31.5 | 1542 |
+| Qwen-Math-Instruct | 9.3 | 4065 | 8.2 | 3590 | 40.5 | 2905 | 78.0 | 1774 | 36.6 | 2888 | 34.0 | 4274 | 34.0 | 3249 |
+| Previous RLVR methods | ||||||||||||||
+| SimpleRL-Zero | 27.0 | - | 6.8 | - | 54.9 | - | 76.0 | - | 34.7 | - | 34.5 | - | 39.0 | - |
 | OpenReasoner-Zero | 16.5 | - | 15.0 | - | 52.1 | - | 82.4 | - | 47.1 | - | 58.7 | - | 45.3 | - |
@@
 | Oat-Zero | 33.4 | - | 11.9 | - | 61.2 | - | 78.0 | - | 43.4 | - | 41.7 | - | 45.9 | - |
-LUFFY | 27.2 | 3120 | 18.1 | 1852 | 61.1 | 1581 | 85.6 | 1370 | 53.6 | 1597 | 52.6 | 1052 | 49.7 | 1762 |
-RL | 22.1 | 1131 | 13.9 | 1527 | 59.5 | 449 | 84.2 | 120 | 44.7 | 433 | 48.3 | 260 | 45.5 | 653 |
-SFT | 26.9 | 16199 | 23.1 | 16609 | 58.9 | 9397 | 85.0 | 3717 | 50.7 | 9769 | 48.6 | 4906 | 48.9 | 10166 |
-ReLIFT | 28.2 | 1666 | 20.1 | 1093 | 64.9 | 943 | 85.2 | 405 | 55.9 | 796 | 52.5 | 1463 | 51.1 | 1061 |
+| Replication of SFT and RL | ||||||||||||||
+| SFT | 26.9 | 7344 | 25.5 | 7059 | 59.8 | 5675 | 85.6 | 3503 | 53.6 | 5662 | 44.4 | 3956 | 49.3 | 5533 |
+| RL | 21.1 | 3287 | 17.5 | 2904 | 62.1 | 2127 | 85.8 | 1277 | 48.7 | 2178 | 46.3 | 1276 | 46.9 | 2175 |
+| Methods combining SFT and RL | ||||||||||||||
+| RL w/ SFT loss | 26.9 | 7344 | 23.1 | 7059 | 59.8 | 5675 | 85.6 | 3503 | 53.6 | 5662 | 44.4 | 3956 | 48.9 | 5508 |
+| SFT then RL (v1) | 29.4 | 5581 | 21.0 | 5328 | 63.1 | 3908 | 87.8 | 2224 | 55.3 | 3772 | 47.6 | 2258 | 50.7 | 3845 |
+| SFT then RL (v2) | 26.8 | 6453 | 23.0 | 6231 | 63.8 | 4586 | 88.4 | 2682 | 54.5 | 4615 | 48.9 | 2638 | 50.9 | 4534 |
+| LUFFY | 27.3 | 5741 | 23.0 | 5267 | 63.5 | 3477 | 85.4 | 2361 | 53.6 | 3758 | 52.6 | 2241 | 50.9 | 3808 |
+| ReLIFT | 28.3 | 5062 | 22.9 | 4571 | 65.1 | 3540 | 87.4 | 2241 | 57.8 | 3352 | 53.9 | 2248 | 52.6 | 3502 |
 4.1 Main Results
-Reasoning Performance.
-Our main results are presented in Table 1, where we compare ReLIFT with several representative zero-RL baselines, as well as direct SFT, on-policy RL, and off-policy RL method LUFFY Yan et al.
-(2025).
-Across five challenging competition-level reasoning benchmarks and one out-of-distribution benchmark, ReLIFT achieves the highest overall accuracy of 51.1%, outperforming all existing baselines and establishing a new state-of-the-art for models based on Qwen2.5-Math-7B.
-Specifically, ReLIFT exceeds the strongest zero RL baseline (Oat-Zero) by a substantial margin of +5.2 points in overall accuracy, and consistently achieves the best or second-best results on all individual benchmarks.
-Furthermore, compared to SFT and RL trained on the same dataset, ReLIFT demonstrates clear superiority: it outperforms SFT by +2.2 points and RL by +5.6 points in overall accuracy.
-ReLIFT also outperforms LUFFY, which combines on-policy and off-policy reasoning traces, despite using only 13% off-policy guidance.
-These results underscore the effectiveness of integrating RL with interleaved fine-tuning for hardest questions, enabling ReLIFT to achieve both higher performance and better generalization.
-Reasoning Efficiency.
-In addition to accuracy, we evaluate the efficiency of different methods by measuring the average response length (LEN) across all benchmarks.
-ReLIFT demonstrates remarkable efficiency, generating shorter solutions (average LEN: 1,061) compared to SFT (average LEN: 10,166), Qwen2.5-Math-Instruct (average LEN: 14,179), and LUFFY, while still achieving superior accuracy.
-This indicates that ReLIFT not only produces more accurate answers but also does so with significantly less computational overhead and more concise reasoning chains.
+| Model | GPU Hours | Demonstrations |
+| SFT | ||
+| RL | ||
+| RL w/ SFT loss | ||
+| SFT then RL v1 | ||
+| SFT then RL v2 | ||
+| LUFFY | ||
+| ReLIFT |
+SOTA performance with Qwen2.5-Math-7B.
+Table 1 presents our main results, comparing ReLIFT against four previous state-of-the-art RLVR methods and replication of SFT and RL, all based on the Qwen2.5-Math-7B model.
+Across five challenging competition-level reasoning benchmarks and one out-of-distribution benchmark, ReLIFT establishes a new state-of-the-art with an overall accuracy of 52.6%, surpassing all baselines.
+Notably, ReLIFT consistently achieves either the best or second-best performance on every individual benchmark, indicating ReLIFT achieves both higher reasoning performance and better generalization.
+Comparing against methods combining SFT and RL.
+When compared with SFT and RL methods trained on the same dataset, ReLIFT outperforms all methods that combine RL and SFT, including RL w/SFT loss, SFT then RL, and LUFFY, as shown in Table 1.
+ReLIFT also produces more concise responses.
+Additionally, an analysis of the required demonstration data and GPU hours in Table 2 reveals that ReLIFT requires less training time and demonstration samples while achieving the best performance.
+These results underscore the effectiveness and efficiency of ReLIFT.
 4.2 Training Dynamics
-Training Dynamics of ReLIFT.
-Figure 3 presents the training dynamics of rewards, response lengths, and the number of hardest questions during RL and ReLIFT training.
-As shown in Figure 3(a), both the RL and ReLIFT models exhibit an increasing trend in rewards, indicating progressive improvement in performance.
-Notably, the reward achieved by ReLIFT consistently surpasses that of standard RL, suggesting that ReLIFT is more effective in optimizing the target objective.
-Regarding response lengths, as depicted in Figure 3(b), RL tends to generate shorter responses as training progresses, which may reflect a diminishing capacity to address more challenging questions.
-In contrast, ReLIFT demonstrates a gradual increase in response length, implying greater computational effort and a higher potential to solve difficult problems.
-Furthermore, when examining the number of hardest questions encountered during training (Figure 3(c)), ReLIFT consistently faces fewer such instances compared to RL, and this number decreases over time.
-Maintaining Exploration.
-As shown in Figure 3(d), the entropy of the RL model steadily declines throughout training, indicating reduced exploration.
+Figure 3 presents the training dynamics of rewards, response lengths, and training entropy during RL and ReLIFT training.
+The reward achieved by ReLIFT consistently surpasses that of RL, suggesting that ReLIFT is more effective in optimizing the target objective.
+Regarding response lengths, RL tends to generate shorter responses as training progresses, which may reflect a diminishing capacity to address challenging questions.
+In contrast, ReLIFT demonstrates a gradual increase in response length, implying a higher potential to solve difficult problems.
+The entropy of the RL model steadily declines throughout training, indicating reduced exploration.
 In contrast, the entropy of ReLIFT remains relatively high and exhibits persistent fluctuations, consistently surpassing that of RL.
 This sustained exploration allows ReLIFT to continue discovering novel solutions, thereby facilitating ongoing performance improvements even in the later stages of training.
-Actually, our method continues to exhibit a upward trend in reward over the final 500 training steps, underscoring the benefits of sustained exploration for long-term optimization.
-3.4 Ablation Study and Analysis
 | Model | AIME-24 | AIME-25 | AMC | MATH-500 | Olympiad | MMLU-Pro | Overall | |||||||
 | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | |
-ReLIFT | 28.2 | 1666 | 20.1 | 1093 | 64.9 | 943 | 85.2 | 405 | 55.9 | 796 | 52.5 | 1463 | 51.1 | 1061 |
-ReLIFT(all) | 7.8 | 26583 | 9.0 | 26704 | 31.6 | 22025 | 54.2 | 16984 | 26.4 | 21670 | 13.7 | 27938 | 23.8 | 23650 |
-ReLIFT(uniform) | 23.2 | 3103 | 19.3 | 2762 | 62.3 | 1389 | 86.0 | 962 | 52.4 | 1426 | 51.4 | 655 | 49.1 | 1716 |
-ReLIFT(random) | 22.1 | 1131 | 13.9 | 1527 | 59.5 | 449 | 84.2 | 120 | 44.7 | 433 | 48.4 | 22644 | 45.5 | 4367 |
-We conduct an ablation study to assess the effectiveness of our ReLIFT approach. Specifically, we compare the following three ablation settings:
-ReLIFT(all): Alternates one fine-tuning step after each RL step on the same batch.
-ReLIFT(uniform): Applies fine-tuning at fixed, uniform intervals (every 8 RL steps), matching the number of demonstrations with ReLIFT.
-ReLIFT(random): Performs interleaved FT.
-The buffer is populated based on the number of Hardest questions, but instead of using the hardest questions, it is randomly filled with other, non-hardest questions.
-As summarized in Table 2, the results show that ReLIFT (all) quickly collapses, probably due to conflicting optimization objectives between RL and FT.
-ReLIFT (uniform) not only reduces accuracy but also leads to longer responses compared to ReLIFT.
-ReLIFT (random) causes a dramatic increase in average response length during training, while still underperforming ReLIFT.
-We also analyze the number of demonstrations of different methods and SFT, as depicted in Figure 4.
-ReLIFT requires only 13% of the demonstrations used in SFT.
-Overall, these findings demonstrate that simply alternating RL and fine-tuning is ineffective; both the scheduling and the selection of fine-tuning training data are critical.
-ReLIFT applies fine-tuning more frequently in the early stages, when the model’s reasoning ability is relatively weak, and focuses on the hardest questions.
-This enables the model to efficiently acquire new knowledge without sacrificing performance or conciseness.
-In contrast, all, uniform, or random fine-tuning scheduling leads to suboptimal results and unnecessary inflation of response length.
+| Qwen2.5-Math-1.5B | ||||||||||||||
+| Qwen-Math-1.5B | 2.8 | 1971 | 1.3 | 2058 | 19.5 | 1533 | 45.4 | 1108 | 18.5 | 1619 | 5.2 | 1982 | 15.5 | 1712 |
+| Qwen-Math-1.5B-Instruct | 10.3 | 4042 | 7.6 | 3764 | 40.0 | 2809 | 77.8 | 1628 | 35.7 | 2834 | 33.7 | 4300 | 34.2 | 3230 |
+| SFT | 12.7 | 7925 | 13.0 | 7719 | 40.9 | 6953 | 71.8 | 5986 | 33.8 | 7231 | 23.5 | 5405 | 32.6 | 6870 |
+| RL | 9.8 | 2425 | 8.5 | 2195 | 44.1 | 1569 | 74.2 | 958 | 37.2 | 1695 | 31.4 | 1075 | 34.2 | 1653 |
+| ReLIFT | 14.3 | 3691 | 10.0 | 3416 | 47.2 | 2207 | 76.4 | 1308 | 39.6 | 2274 | 31.7 | 1724 | 36.5 | 2437 |
+| Qwen2.5-7B | ||||||||||||||
+| Qwen-7B | 6.2 | 1632 | 2.4 | 1370 | 31.1 | 1076 | 63.8 | 617 | 26.5 | 1068 | 32.8 | 945 | 27.1 | 1118 |
+| Qwen-7B-Instruct | 11.5 | 1939 | 6.1 | 1500 | 41.0 | 1629 | 73.0 | 1753 | 37.3 | 1550 | 53.1 | 3325 | 37.0 | 1949 |
+| SFT | 15.7 | 7786 | 18.6 | 7533 | 49.8 | 6424 | 80.8 | 4252 | 44.3 | 6234 | 54.6 | 4869 | 44.0 | 6183 |
+| RL | 15.5 | 1784 | 13.0 | 1487 | 50.4 | 1422 | 80.0 | 917 | 42.2 | 1438 | 57.6 | 954 | 43.1 | 1334 |
+| ReLIFT | 19.1 | 5522 | 14.7 | 4831 | 51.9 | 3792 | 81.6 | 2211 | 46.1 | 3649 | 56.4 | 2173 | 45.0 | 3696 |
+| LLaMA-3.1-8B | ||||||||||||||
+| LLaMa-8B-Instruct | 5.6 | 1396 | 0.7 | 1401 | 20.7 | 1012 | 44.0 | 622 | 13.8 | 1063 | 38.1 | 357 | 20.5 | 975 |
+| SFT | 0.8 | 2033 | 1.5 | 2032 | 11.5 | 2007 | 28.6 | 1903 | 8.6 | 1997 | 28.2 | 1558 | 13.2 | 1922 |
+| RL | 1.8 | 940 | 0.0 | 909 | 10.8 | 797 | 28.2 | 616 | 7.3 | 790 | 39.4 | 549 | 14.6 | 767 |
+| ReLIFT | 1.3 | 1236 | 0.2 | 1272 | 11.9 | 1048 | 35.2 | 741 | 11.0 | 1050 | 44.2 | 650 | 17.3 | 1000 |
+4.3 Ablation Study and Analysis
+We conduct an ablation study to understand what makes ReLIFT effective.
+Our findings reveal that the scheduling of fine-tuning and the selection of training data are crucial.
+Specifically, we compare the following three ablation settings: ReLIFT(all), which alternates one fine-tuning step after each RL step on the same batch, with the number of demonstrations matching SFT; ReLIFT(uniform), which applies fine-tuning at fixed, uniform intervals (every 8 RL steps), with the number of demonstrations matching ReLIFT; ReLIFT(random) which performs interleaved fine-tuning.
+The buffer is populated based on the number of Hardest questions, but instead of using the hardest questions, it is randomly filled with non-hardest questions.
+As shown in Figure 5, ReLIFT (all) quickly collapses, probably due to conflicting optimization objectives between RL and FT.
+Similarly, using a uniform or random schedule for FT, as in ReLIFT(uniform) and ReLIFT(random), leads to lower accuracy and longer responses.
+In contrast, ReLIFT strategically applies fine-tuning more frequently in early training stages and focuses on the most challenging questions.
+This approach allows the model to efficiently gain new knowledge without sacrificing performance or conciseness.
+More detailed results are in Appendix D.
+We also study the effect of the entropy coefficient, , on the fine-tuning step, using Qwen2.5-Math-7B.
+As seen in Figure 5, an of 0.0001 prove optimal, achieving the highest accuracy of 52.6.
+Varying in either direction from this value significantly decreased performance.
+These results highlight the importance of the entropy loss term for effectively blending SFT with RL.
+Therefore, we use consistently in our experiments.
 4.4 Extension to More Models
-Model | AIME-24 | AIME-25 | AMC | MATH-500 | Olympiad | MMLU-Pro | Overall | |||||||
-ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | |
-Qwen-Math-1.5B | 3.0 | 4105 | 1.4 | 4291 | 19.4 | 3480 | 44.4 | 2637 | 16.7 | 3487 | 5.0 | 6260 | 15.0 | 4043 |
-Qwen-Math-1.5B-Instruct | 8.1 | 5722 | 6.6 | 4468 | 36.9 | 4614 | 66.0 | 6148 | 30.7 | 5526 | 24.2 | 8360 | 28.8 | 5806 |
-SFT | 12.7 | 18060 | 13.0 | 19271 | 40.9 | 12538 | 71.8 | 5879 | 33.8 | 12906 | 23.5 | 11767 | 32.6 | 13403 |
-RL | 10.2 | 1487 | 8.7 | 1187 | 44.7 | 796 | 71.4 | 423 | 34.5 | 832 | 28.7 | 756 | 33.0 | 913 |
-ReLIFT | 13.1 | 6294 | 8.8 | 5866 | 42.5 | 3856 | 73.6 | 2750 | 36.0 | 3937 | 30.1 | 3969 | 34.0 | 4421 |
-Qwen-7B | 5.1 | 3146 | 1.4 | 3518 | 22.6 | 2611 | 37.2 | 1811 | 19.0 | 2429 | 39.8 | 434 | 20.9 | 2324 |
-Qwen-7B-Instruct | 10.3 | 1299 | 6.5 | 562 | 34.9 | 1043 | 67.6 | 605 | 29.8 | 907 | 53.3 | 7413 | 33.7 | 1804 |
-SFT | 15.7 | 19363 | 18.6 | 18070 | 49.8 | 11349 | 80.8 | 4453 | 44.3 | 11528 | 54.6 | 5188 | 43.9 | 11658 |
-RL | 13.3 | 4371 | 4.2 | 3988 | 40.2 | 2845 | 65.4 | 1971 | 31.4 | 2852 | 56.6 | 180 | 35.2 | 2867 |
-ReLIFT | 19.1 | 12197 | 14.7 | 9284 | 51.9 | 5909 | 81.6 | 2395 | 46.1 | 6160 | 56.4 | 2550 | 45.0 | 6415 |
-To evaluate the general applicability of our method, we extend ReLIFT to both a smaller base model, Qwen2.5-Math-1.5B Yang et al.
-(2024), and a weaker base model, Qwen2.5-7B Team (2024).
-Table 3 presents the performance across five challenging competition-level math benchmarks and one out-of-domain benchmark.
-On both models, ReLIFT consistently achieves substantial improvements, surpassing both SFT and RL.
-Across all benchmarks, our method ranks first or second, while maintaining an acceptable response length.
-These results underscore the applicability of ReLIFT to achieve both higher performance and better generalization.
+To evaluate the generalizability of our method, we extend ReLIFT to three additional base models: the smaller Qwen2.5-Math-1.5B Yang et al.
+(2024), the weaker Qwen2.5-7B Team (2024), and Llama3.1-8b Dubey et al.
+(2024), which has a different architecture.
+As shown in Table 3, ReLIFT consistently achieves substantial improvements across all base models and six benchmarks, surpassing both SFT and RL.
+Our method consistently achieves a top-two ranking while maintaining an acceptable response length.
+These results underscore the broad applicability of ReLIFT in achieving superior performance and generalization across diverse model scales and architectures.
 5 Related Works
 5.1 RL-Driven Emergence of Reasoning Abilities
-Reinforcement learning has recently emerged as a promising approach for enhancing the reasoning capabilities of large language models (LLMs).
-Early breakthroughs, such as O1 Jaech et al.
-(2024), DeepSeek R1 Guo et al.
-(2025), and Kimi 1.5 Team et al.
-(2025), have demonstrated that applying Reinforcement Learning with Verifiable Rewards (RLVR) can lead to significant improvements in complex reasoning tasks.
-Notably, DeepSeek R1 further shows that applying RLVR directly to a base model—without the need for intermediate supervised fine-tuning—can also effectively enhance the reasoning abilities of LLMs.
+RL has recently emerged as a promising approach for enhancing the reasoning capabilities of LLMs.
+Early breakthroughs, such as OpenAI o-series Jaech et al.
+(2024); OpenAI (2025), DeepSeek-R1 Guo et al.
+(2025), and Kimi k-series Team et al.
+(2025b; a), have demonstrated that applying RLVR can lead to significant improvements in complex reasoning tasks.
+Notably, DeepSeek-R1 further shows that applying RLVR directly to a base model, without the need for intermediate SFT.
 Building on these foundations, Light-r1 Xie et al.
-(2025) incorporated curriculum learning to further refine reasoning skills, while Logic-rl Xie et al.
-(2025) adopted logic-driven reward functions to improve general reasoning ability.
+(2025) incorporates curriculum learning to refine LLM’s reasoning skills, while Logic-rl Xie et al.
+(2025) adopts logic-driven reward functions to improve general reasoning ability.
 Deepscaler Luo et al.
-(2025) introduced Iterative Context Lengthening, a method that trains models with progressively longer contexts as their performance improves, thereby reducing both monetary costs and end-to-end training time.
-In addition, Oat-Zero Liu et al.
-(2025) was trained from Qwen2.5-Math-7B using rule-based rewards, and proposed removing the standard deviation in GRPO advantage computation as well as token-level normalization in policy loss computation.
-PRIME-Zero Cui et al.
-(2025) leveraged policy rollouts and outcome labels, utilizing implicit process rewards to further enhance model performance.
+(2025) trains LLMs with progressively longer contexts as their performance improves.
+Based on GRPO, DAPO Yu et al.
+(2025) uses clip-higher to prevent entropy collapse, dynamic sampling for improved efficiency, and token-level policy gradient loss with overlong reward shaping to stabilize training.
 5.2 Training Paradigms
-Despite significant advances in reinforcement learning (RL), Yue et al.
-Yue et al.
-(2025) argue that RL with value-based rewards (RLVR) does not enable large language models (LLMs) to acquire genuinely new reasoning abilities beyond those present in their base models.
-Instead, RL primarily enhances performance by biasing the model toward reasoning paths it already knows that are more likely to yield rewards.
-In other words, most RL methods optimize within the model’s existing knowledge rather than facilitating the acquisition of new information, thereby limiting the LLM’s overall capabilities.
-In contrast, directly fine-tuning on high-quality demonstration data can introduce new knowledge and abilities to the model, and this approach is widely used to adapt LLMs to downstream reasoning tasks.
-However, supervised fine-tuning (SFT) requires high-quality, detailed demonstration data and tends to encourage memorization of the training data, often resulting in poor generalization to out-of-distribution scenarios Chu et al.
-(2025).
-Some approaches attempt to combine supervised fine-tuning (SFT) and reinforcement learning (RL).
-Typically, these methods first use SFT to stabilize the model’s output format, then apply RL to further improve performance and potentially unlock more advanced reasoning abilities in LLMs Guo et al.
+Despite these significant advances, Yue et al.
+(2025) argue that RLVR does not enable LLMs to acquire genuinely new reasoning abilities beyond those present in their base models, as RLVR-trained models perform worse on pass@k than the base models at higher k values, indicating limited reasoning capability expansion.
+Similarly, Zhao et al.
+(2025) demonstrate that RL creates an ”echo chamber” effect, amplifying specific behavioral patterns already present in the pre-training data while suppressing others, thereby reinforcing existing behaviors rather than creating new reasoning capabilities.
+Furthermore, Cheng et al.
+(2025) finds that RL stifles exploration, causing models to converge on narrow behaviors and leading to performance plateaus on complex reasoning tasks.
+Research suggests that current RLVR methods do not fully leverage the potential of RL to create genuinely novel reasoning abilities in LLMs.
+In contrast, directly fine-tuning on demonstration data introduces new knowledge to LLMs.
+However, SFT requires high-quality, detailed demonstration data and tends to encourage memorization of the training data, often resulting in poor generalization to out-of-distribution scenarios Chu et al.
+(2025); Chen et al.
+(2025); Team et al.
+(2025b).
+Some approaches attempt to combine SFT and RL.
+Typically, these methods first use SFT to stabilize the LLM’s ability and then apply RL to unlock more advanced reasoning skills in LLMs Guo et al.
 (2025); Wen et al.
 (2025).
-However, such combined methods remain limited: SFT is highly dependent on the quality of demonstration data, and the subsequent RL phase is still confined to the knowledge already present in the SFT-trained model, making it difficult for the model to acquire genuinely new information.
+However, such combined methods remain limited: SFT depends significantly on the quality of demonstration data, and the subsequent RL phase is still confined to the knowledge present in the SFT-trained model, making it difficult for the model to acquire genuinely new information.
 LUFFY Yan et al.
 (2025) seeks to address these limitations by augmenting RL with off-policy reasoning traces, dynamically balancing imitation and exploration through a combination of off-policy demonstrations and on-policy rollouts during training.
-However, this approach requires the preparation of off-policy data in advance and relies heavily on probability-sharpening techniques, which may restrict its applicability.
-In contrast, our work proposes a simpler and more direct method that combines RL and SFT to leverage the strengths of both paradigms.
-Our approach encourages LLMs to explore when possible and to incorporate new information as needed, thereby pushing RL-driven reasoning beyond the boundaries of existing knowledge.
+Nevertheless, this approach necessitates the preparation of off-policy data in advance and relies heavily on probability-sharpening techniques, which may restrict its applicability.
+In contrast, ReLIFT proposes a simpler and more direct method that integrates RL and SFT to leverage the strengths of both paradigms.
+Our approach encourages LLMs to explore when possible and integrate new information as needed, thereby pushing RL-driven reasoning beyond the boundaries of existing knowledge.
 6 Conclusion
 In summary, we systematically analyze the respective strengths of RL and SFT for LLM reasoning, finding that RL is more effective for learning easier questions while SFT is crucial for solving challenging questions.
-Building on these insights, we introduce ReLIFT, an approach that interleaves RL with online Fine-Tuning on the hardest questions, achieving state-of-the-art results with significantly less demonstration data and more concise solutions.
+Building on these insights, we introduce ReLIFT, an approach that interleaves RL with online fine-tuning on the hardest questions.
+ReLIFT achieves state-of-the-art results with significantly less demonstration data, reduced training time, and more concise solutions.
 Future work will focus on scaling ReLIFT to larger models and developing more effective strategies for coordinating SFT and RL to further advance LLM reasoning and generalization.
-Appendix A Experiment Details
-A.1 Main Experiment Details
-Training.
-In addition to Qwen2.5-Math-7B, we extend ReLIFT to both Qwen2.5-Math-1.5B Yang et al.
-(2024) and Qwen2.5-7B Team (2024).
-Since the maximum sequence length for Qwen2.5-Math-1.5B is 4096, which is insufficient for our tasks, we follow the approach used for Qwen2.5-Math-7B by increasing the RoPE theta from 10,000 to 40,000 and expanding the window size to 16,384.
-To ensure fairness, we maintain 8 rollouts per prompt for all RL-trained models.
-The learning rate is fixed at .
-All training experiments are conducted on 2 8 A800 GPUs.
-We train all RL models for 500 steps and all SFT models for three epochs.
-For all experiments, we use verl Sheng et al.
+Appendix A The Use of Large Language Models
+We use LLMs to polish our writing. They act as a sophisticated assistant, improving the clarity, grammar, and style of our text.
+Appendix B Main Experiment Details
+RL Implementation.
+For RL, we omit the KL divergence loss term following Luo et al.
+(2025); Yan et al.
 (2025).
+Consistent with Dr.GRPO Liu et al.
+(2025) and LUFFY Yan et al.
+(2025), we exclude both length normalization and standard error normalization from the GRPO loss 1 in all experiments.
+The rollout batch size is set to 128, the number of rollout is 8, the update batch size is 64, and the learning rate is .
+A temperature of 1.0 is used during rollout generation for exploration.
+For Qwen2.5-Math-7B, We train 600 steps for all methods except SFT then RL, to be near with the GPU hour of ReLIFT.
+For other base models, we train 500 steps.
+All experiments are conducted on 8 x A800 GPUs.
+SFT Implementation.
+For all our SFT models, we utilize the OpenR1-Math-46k-8192 dataset.
+Following a hyperparameter search on a sampled small-scale dataset, we evaluate batch sizes of 64, 128, and 256, in conjunction with learning rates of , and .
+The optimal configuration identified includes a batch size of 64 and a learning rate of .
+Employing the LLaMA-Factory framework Zheng et al.
+(2024), we train our models for 3 epochs.
+RL w/ SFT Loss Implementation. We compute the on-policy loss on 7 on-policy samples and SFT loss on 1 off-policy sample per prompt. The remaining setup is consistent with our RL experiments.
+SFT then RL Implementation.
+We conduct SFT using the same settings outlined in the SFT Implementation and train for 3 epochs accordingly.
+Subsequently, the RL phase follows the same configuration as described in RL Implementation, and we train the RL phase for 300 steps.
+ReLIFT Implementation.
+During interleaved fine-tuning, we make use of demonstrations collected beforehand.
+The RL and fine-tuning configurations for ReLIFT are consistent with the settings previously described except learning rate.
+For both the RL and interleaved SFT phases, the learning rates are set to .
 Evaluation.
@@
 Results are verified using math-verify Face (2024).
-For the baselines, we independently reproduce and verify the results of the LUFFY Yan et al.
-(2025), whereas the results for the other zero-RL models are reported as in the LUFFY paper.
-Chat Template.
+For the baselines, we independently reproduce and verify the results of all methods combining SFT and RL, such as LUFFY Yan et al.
+(2025), whereas the results for previous RLVR methods are reported as in the LUFFY paper.
+Reward Function. To evaluate the impact of our method, we adopt a simple reward function as below. All training experiments employ the same reward function.
+Qwen2.5-Math Models.
+All evaluations are conducted using vllm Kwon et al.
+(2023), with a temperature setting of 0.6 and a maximum token limit of 8192.
+Results are validated using math-verify Face (2024).
+For baseline methods, we independently reproduce and verify results for all approaches combining SFT and RL, including LUFFY Yan et al.
+(2025).
+The results for other RL models are reported following the findings presented in the LUFFY paper.
+Llama-3.1-8B.
+Because the full OpenR1-Math-46k-8192 dataset proved too challenging for Llama-3.1-8B, we created a smaller, more manageable dataset.
+We selected 11,000 samples from OpenR1 where DeepSeek-R1 provided the correct solution and the solution length was under 2,048 tokens.
+The SFT was then performed using a learning rate of , as suggested by Liu et al.
+(2025).
+The maximum response length is set to 2048 tokens for RL training and evaluation.
+Motivation Experiment Details.
+We randomly sample a subset of 8,000 examples from the training dataset described in Section 3 as the training set, and select 1,000 examples as the validation set.
+RL and SFT are both conducted on this subset dataset for 120 steps, using the same parameters as in our main experiments.
+Appendix C Chat Template
 To encourage the base model to reason more effectively, following Yan et al.
 (2025), we adopt a complex chat template as shown in Figure 6.
-All training experiments and evaluation utilize the same chat template.
-Reward Function. To evaluate the impact of our method, we adopt a simple reward function as below. All training experiments employ the same reward function.
-A.2 Motivation Experiment Details
-Dataset.
-We randomly sample a subset of 8,000 examples from the training dataset described in 3.1 as the training set, and select 1,000 examples as the validation set.
-RL and SFT are both conducted on this subset dataset for 120 steps, using the same parameters as in our main experiments.
+All our trained models, except LLaMA-3.1-8B, share the same system prompt for training and inference.
+For LLaMA-3.1-8B, we do not use the above system prompt as we find the model cannot follow such an instruction. Thus, we use a simplified version as shown in Figure 7.
+Appendix D Detailed Experiment Results
+| Model | AIME-24 | AIME-25 | AMC | MATH-500 | Olympiad | MMLU-Pro | Overall | |||||||
+| ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | ACC | LEN | |
+| ReLIFT | 28.3 | 5062 | 22.9 | 4571 | 65.1 | 3540 | 87.4 | 2241 | 57.8 | 3352 | 53.9 | 2248 | 52.6 | 3502 |
+| ReLIFT(all) | 7.8 | 7647 | 9.0 | 7648 | 31.6 | 6653 | 54.2 | 5320 | 26.4 | 6600 | 13.7 | 6590 | 23.8 | 6743 |
+| ReLIFT(uniform) | 23.2 | 6113 | 19.3 | 5318 | 62.3 | 4016 | 86.0 | 1971 | 52.4 | 2926 | 51.4 | 1952 | 49.1 | 3716 |
+| ReLIFT(random) | 22.1 | 4616 | 13.9 | 4676 | 59.5 | 5271 | 84.2 | 5820 | 44.7 | 5187 | 48.4 | 6020 | 45.5 | 5268 |
+The comprehensive results for ReLIFT(all), ReLIFT(uniform), and ReLIFT(random) are presented in Table 4.
+These findings indicate that merely alternating between RL and fine-tuning is insufficient; both the scheduling and the selection of fine-tuning training data are crucial components for success.
+Appendix E Reasoning Behaivers
+To analyze the distinct behaviors of the RL, SFT, and ReLIFT models, we first curate a pool of keywords corresponding to three key cognitive actions: summarizing, rethinking, and planning.
+We then quantified the frequency of these keywords within the AIME25 responses generated by each model.
+To ensure a fair comparison across varying response lengths, all keyword counts are normalized by the total number of tokens.
+As shown in Figure 8, our analysis reveals several key differences in how the models approach problem-solving.
+Summarizing: The SFT model exhibits a notably higher usage of summary keywords, suggesting it often prioritizes synthesizing and restating information.
+The ReLIFT model uses slightly more summary keywords than the RL model, but both are far behind SFT.
+Rethinking: In contrast, the ReLIFT model uses significantly fewer keywords related to rethinking or re-evaluation.
+This finding suggests a higher level of efficiency and confidence in its initial approach, requiring less internal correction or revision during the problem-solving process.
+Planning: The most striking difference is in the use of planning keywords.
+The ReLIFT model demonstrates a substantial increase in these keywords, indicating a more robust and sophisticated approach to strategic planning.
+This suggests that ReLIFT is not just more efficient but also more capable of developing a logical and powerful problem-solving trajectory.
+Appendix F Case Study
+AIME25 I Problem 3: The first step to solve this problem is to find three possible combinations , and .
+The RL response fails to include one possible combination, while the SFT response requires excessive analysis to obtain the three combinations.
+In contrast, the ReLIFT response is more concise and obtains all three combinations.